# High-Speed Event Vision-Based Tactile Roller Sensor for Large Surface Measurements

### v1

@@ Abstract
Inspecting large-scale industrial surfaces like aircraft fuselages for quality control requires capturing their precise 3D surface geometry at high resolution. Vision-based tactile sensors (VBTSs) offer high local resolution but require slow ’press-and-lift’ measurements stitched for large areas. Approaches with sliding or roller/belt VBTS designs provide measurements continuity. However, they face significant challenges respectively: sliding struggles with friction/wear and both approaches are speed-limited by conventional camera frame rates and motion blur, making large-area scanning time consuming. Thus, a rapid, continuous, high-resolution method is needed. We introduce a novel tactile sensor integrating a neuromorphic camera in a rolling mechanism to achieve this. Leveraging its high temporal resolution and robustness to motion blur, our system uses a modified event-based multi-view stereo approach for 3D reconstruction. We demonstrate state-of-the-art scanning speeds up to 0.5 m/s, achieving Mean Absolute Error below 100 microns – 11 times faster than prior continuous tactile sensing methods. A multi-reference Bayesian fusion strategy enhances accuracy (reducing MAE by 25.2% compared to EMVS) and mitigates curvature errors. We also validate high-speed feature recognition via Braille reading 2.6 times faster than previous approaches.
@@ I Introduction
Surface metrology and surface inspection are crucial elements in quality assurance across diverse industries, particularly aerospace and automotive manufacturing. Precise inspection is required to identify characteristics like paint quality, coating integrity, and subtle defects such as cracks, nicks, and dents [1, 2, 3]. Often, achieving a resolution of 0.1 mm or lower is necessary to accurately classify these features and ensure component integrity and safety [4]. Traditional contact-based methods, including high-precision profilometers [5, 6] or microscopic techniques [7, 8, 9], offer high resolution locally but become exceedingly time-consuming when applied to large surface areas due to their sequential, point-by-point or small-patch measurement nature. Non-contact optical methods, such as cameras, laser scanners, or structured light systems [2, 10, 11, 12, 13, 14], can significantly accelerate inspection by capturing data over wider areas. However, these methods often lack robustness; their performance can be compromised by variations in ambient lighting, motion blur when attempting high-speed scanning, or challenging surface optical properties like high reflectivity or transparency [15]. This highlights a critical need for sensing modalities that combine high resolution and robustness with the capability for rapid inspection of large, potentially complex surfaces.
Vision-Based Tactile Sensors (VBTSs), particularly those derived from the GelSight concept [3, 16, 17], have emerged as a powerful contact-based alternative offering a unique blend of high-resolution geometric and textural sensing. These sensors typically employ a camera to view the deformation of a soft, reflective elastomer membrane pressed against a target surface. This principle allows for detailed surface capture, often with inherent robustness to external lighting conditions due to internal illumination [3, 18]. Numerous variations of GelSight have been developed, enhancing capabilities for robotic manipulation, complex shape capture, and integration into compact form factors [19, 20, 21, 22].
Despite their advantages for detailed local analysis, the standard planar configuration of most VBTSs necessitates a discrete ”press-and-lift” interaction mode. This process, where the sensor is repeatedly pressed, lifted, and moved, becomes prohibitively time-consuming when inspecting large surface areas, requiring numerous measurements to be stitched together [4], thereby limiting their use for efficient large-scale tactile measurement.
@@
To enable continuous tactile measurement, various designs moving beyond the static planar configuration have emerged. Attempts at direct sliding of planar sensors face significant friction, wear, and motion blur challenges [23]. Roller-based [24, 25] and belt-based [26] VBTS designs replace the static contact with a dynamic rolling interface, allowing continuous data acquisition. TouchRoller, for example, demonstrated 2D texture mapping over moderate areas [24], while GelBelt provides a larger contact patch via its belt mechanism [26]. Despite these mechanical advancements for measurement continuity, these designs face limitations. Reconstructing accurate 3D geometry from the depth-variant information gathered by cylindrical rollers is complex, primarily because the contact depth is not uniform across the curved sensing area. This variation distorts the captured tactile image and complicates the mapping between observed elastomer deformation and the actual 3D surface geometry [27]. More critically, the operational speed of systems relying on conventional cameras is constrained. Frame rate limitations and motion blur become significant issues as scanning speed increases. This problem is particularly acute for VBTS, which often require close-up imaging with a shallow depth of field to resolve fine surface details; any slight vibration during the camera’s exposure time at high speeds drastically worsens motion blur, degrading reconstruction accuracy. Consequently, current roller/belt tactile sensors are restricted to relatively low velocities (e.g., typically well below 0.1 m/s for 3D reconstruction [24, 26]), insufficient for truly high-throughput industrial inspection of large parts.
@@
Neuromorphic vision sensors, or event cameras, provide a promising technological solution to overcome these speed barriers [28, 29]. Unlike frame-based cameras, event cameras feature pixels that respond independently and asynchronously to logarithmic brightness changes, generating data with microsecond-level temporal resolution, high dynamic range, and inherent resistance to motion blur [30, 31, 32]. These characteristics are ideally suited for capturing the rapid, dynamic deformations occurring at the contact interface of a high-speed rolling tactile sensor. Crucially, integrating event cameras into tactile roller designs holds considerable potential for achieving high-resolution continuous tactile sensing at higher speeds with significantly lower computational power, particularly for surfaces with sparse details, compared to conventional VBTSs. While Neuromorphic VBTSs (NVBTS) have been explored for applications like robotic grasping and slip detection (e.g., Evetac [33]) or contact analysis [34, 35, 36, 37], their potential for enabling high-speed, continuous 3D large surface reconstruction using a roller mechanism has not been realized. A significant gap exists for a tactile sensing system capable of combining the continuous scanning geometry of a roller with the high temporal fidelity and motion robustness of event-based vision for large-scale surface measurement.
@@
This paper introduces, to the best of our knowledge, the first high-speed neuromorphic vision-based roller tactile sensor specifically designed for large surface measurements (illustrated in Figure 1). Our primary contributions are:
@@
    We present a novel rolling neuromorphic vision-based tactile sensor for 3D reconstruction and feature recognition of surfaces at speeds reaching up to 0.50.50.5 m/s, 11 times faster than the previous approaches.
@@
    We propose a modified event-based multi-view Stereo (EMVS) algorithm. Our approach enhances depth estimation and consistency on curved tactile sensors by fusing depth maps from multiple reference views. Furthermore, we include a corresponding calibration routine to optimize its parameters.
  4. 4.
    We rigorously test the proposed roller-based tactile sensor (RBTS) in 3D reconstruction tasks and compare its performance against the state-of-the-art for large surface inspection with regards to speed and accuracy.
@@ II Related Works
Reconstructing a surface’s three-dimensional geometry is crucial in many applications, from automated manufacturing to quality control. Mechanical profilometers have long been used for high-precision surface measurements, meticulously tracing the surface point-by-point. However, this inherent sequential nature makes them slow, especially when dealing with complex geometries or large areas [6]. Non-contact optical techniques, like laser scanning and structured light, offer a significant speed advantage by rapidly capturing surface data over a wider area [38, 39]. These methods boast resolutions around 0.1 mm [16], but their reliance on light makes them susceptible to issues with highly reflective or transparent materials, as well as variations in ambient lighting conditions [15]. Furthermore, certain laser scanning systems, particularly higher-power ones used for speed or range, can introduce eye-safety concerns requiring adherence to safety standards and protocols [40].
@@
For applications demanding the utmost precision, techniques like white light interferometry, confocal microscopy, and atomic force microscopy (AFM) are utilized [7, 8, 9]. These methods can resolve surface details down to the nanometer scale, revealing intricate textures and features. However, they often involve complex setups, lengthy measurement times, and significant expense, limiting their practical use for large-scale inspection tasks.
II-B Vision-Based Tactile Sensors
@@
Vision-Based Tactile Sensors (VBTSs) have emerged as compelling alternatives for surface sensing, offering robustness, versatility, and ease of use. The GelSight sensor pioneered this field, making use of a soft, deformable surface with camera-based observation to extract detailed surface information utilizing photometric stereo [16]. By analyzing the deformation of the elastomer, typically coated with a reflective layer, VBTSs can capture high-resolution data on surface geometry and texture, even under varying lighting [3, 18]. From the original GelSight concept, a diverse range of VBTS designs have emerged to enhance performance and broaden application: GelSlim for more compact integrations [19], GelSight Fin Ray for soft compliant grippers [20], and GelSight Svelte for human shape mimicking [21] to name a few. Crucially, simulation tools have evolved in parallel, accelerating the design and deployment of novel VBTS solutions [41, 42, 43, 44].
@@
However, conventional VBTSs often face challenges when it comes to continuous surface scanning. The rigid backing of the elastomer limits its ability to slide smoothly across the surface, leading to distortions and potential damage to the sensing membrane. Consequently, these sensors often require a discrete ”press-and-lift” approach, making them ill-suited for rapid inspection of large areas; for example, the commercial GelSight Mobile™ Series 2 requires approximately 0.1 seconds of exposure time per measurement patch [45], rendering large area coverage time-consuming [4]. An attempt to address this limitation using continuous sliding with a conventional VBTS (DIGIT) was explored by Potdar et al. for high-speed Braille reading [23]. While demonstrating the potential for sliding interactions, their approach, relying on a conventional camera, encountered significant motion blur artefacts at higher velocities, leading to a notable decrease in recognition accuracy as the sliding speed increased. Furthermore, the inherent friction during sliding poses a significant challenge for conventional VBTS elastomers, potentially causing rapid wear or damage to the delicate sensing surface, an issue Potdar et al. mitigated partially through the use of protective tape [23]. While reducing wear, such protective layers can introduce optical diffusion, potentially limiting the resolution of very fine textures, and may necessitate more precise control over maintaining perpendicular contact during sliding to ensure consistent data acquisition.
II-C Neuromorphic Vision-Based Tactile Sensing
@@
Neuromorphic vision sensors, or event cameras, offer a transformative technology for high-speed tactile sensing by directly addressing the limitations of conventional cameras [28]. Their asynchronous pixel operation leads to microsecond temporal resolution, high dynamic range, low latency, and inherent suppression of motion blur [30, 31, 32]. These properties are exceptionally well-suited to capturing the rapid, dynamic changes in elastomer deformation during high-speed continuous tactile scanning. Several Neuromorphic VBTSs (NVBTS) have leveraged these advantages. Initial explorations included using Gaussian processes on event data for contact force estimation [46] or employing spiking neural networks (SNNs) for texture classification (NeuroTac [34]), though SNNs can be challenging to train for regression tasks like pose or force prediction. More recent systems like Evetac [33] utilize event-based marker tracking for robust slip detection during robotic grasping, demonstrating the utility of events for dynamic contact analysis. Graph neural networks have been applied to event streams for contact angle prediction (Tactigraph [35]), and other works have explored generating image-like representations from events for proprioception or other analyses [36]. While these studies validate the potential of NVBTS, their focus has predominantly been on grasping-related tasks (slip, force, contact classification) or local texture analysis. Moreover, for these use cases, most NVBTS suffer from decreased spatial resolution due to the reliance on markers, which is detrimental for surface reconstruction and depth estimation [33].
II-D Roller-Based and Large Surface Tactile Sensors
@@
Roller-based tactile sensors represent a promising avenue for achieving continuous tactile sensing, particularly for large surfaces where traditional planar sensors are inefficient. Early work explored cylindrical designs for specific detection tasks, such as the sensor developed by Shimonomura et al. [25] for detecting foreign bodies like food impurities during rolling contact. Similarly, studies in robotic in-hand manipulation have successfully utilized a roller grasper, often equipped with high-resolution tactile feedback, for tasks like object insertion and regrasping [47, 48], demonstrating the utility of rolling contact in dynamic scenarios beyond simple surface mapping.
@@
Transitioning towards surface property characterization, TouchRoller [24] demonstrated the potential for mapping surface textures by continuously rolling a sensor element across a fabric sample. It achieved coverage of an 8cm $\times$ 11cm area in 10 seconds, yielding a 2D texture map. Notably, its operational speed was limited to approximately 11 mm/s; the authors noted that higher speeds resulted in motion blur within the conventional camera frames, hindering performance. This motion blur problem is particularly acute for VBTSs, which require close-up imaging to resolve fine surface details. Operating at such close proximity means that even moderate rolling speeds translate into high velocities of features across the image sensor plane; during the camera’s finite exposure time, these features traverse a larger number of pixels, resulting in significant motion blur compared to imaging from a distance. Furthermore, TouchRoller did not perform 3D surface reconstruction. Addressing 3D reconstruction, GelRoller [49] employed a cylindrical design integrating photometric stereo techniques for local surface normal estimation from single frames. For reconstructing larger surfaces, they proposed stitching multiple local reconstructions using Iterative Closest Point (ICP) registration. While demonstrating 3D capability, this work did not explicitly analyze or optimize the operational speed of the rolling process.
@@
A fundamental challenge inherent to cylindrical roller designs, however, is that the tactile information gathered varies significantly with contact depth due to the sensor’s curvature. This leads to inconsistencies and potential distortions in the acquired tactile images, complicating accurate reconstruction. Researchers have explored methods to mitigate this depth inconsistency. Li et al. [27] proposed a cyclic image fusion strategy, processing information from different contact phases within the rolling sequence using wavelet transforms to enhance information extraction and reduce bias from the curved contact.
@@
An alternative approach addresses the issue mechanically: the GelBelt sensor [26] utilizes a flat elastomeric belt moving between two wheels, creating a larger, more uniform contact area akin to a conveyor. This design avoids the depth variation inherent in direct cylindrical contact and enabled continuous sensing at higher reported speeds, up to 45 mm/s. However, this design comes with some drawbacks. It relies on sliding motion to achieve high-speed operation, using an air gap and a smooth surface to reduce friction. While this allows GelBelt to move quickly, it creates challenges for lighting and can lead to wear over time, which may affect the system’s accuracy and repeatability. The design also struggles with curved surfaces. Its large, flat structure makes it hard to access concave areas, and it doesn’t perform well on convex surfaces either. In this case, the flat sensor doesn’t match the shape of the curved surface, leading to a limited sensing area, similar to problems seen in roller-based systems.
@@
Furthermore, despite these advances in measurement continuity, including GelBelt’s increased speed, systems relying on conventional cameras remain fundamentally constrained by the trade-off between speed and accuracy due to motion blur. As demonstrated by GelBelt’s own results showing decreased accuracy at higher speeds [26], pushing scanning velocities further significantly degrades the quality of the captured tactile information and subsequent reconstructions. While naively increasing camera frame rates might seem like a straightforward solution to combat motion blur, it introduces substantial challenges regarding the increased computational complexity required for real-time processing, significantly larger data storage requirements, and higher energy consumption, potentially limiting practical deployment. In turn, the potential for using event-based vision in roller sensors to achieve similar resolution at low computational power and significantly higher speeds for continuous tactile sensing is considerable. However, this capability has yet to be demonstrated, particularly for demanding large-scale inspection tasks.
III Methodology
III-A Mechanical Design and Fabrication

### v2

# They See Me Rolling: High-Speed Event Vision-Based Tactile Roller Sensor for Large Surface Inspection

@@ Abstract
Inspecting large-scale industrial surfaces like aircraft fuselages for quality control requires precise, high-resolution 3D geometry. Vision-based tactile sensors (VBTSs) offer high local resolution but require slow ’press-and-lift’ measurements for large areas. Sliding or roller/belt VBTS designs provide continuous measurement but face significant challenges: sliding suffers from friction/wear, while both are speed-limited by camera frame rates and motion blur. Thus, a rapid, continuous, high-resolution method is needed. We introduce a novel neuromorphic tactile roller sensor. It uses a modified event-based multi-view stereo algorithm for 3D reconstruction, leveraging high temporal resolution and motion blur robustness. This reconstruction is most effective for surfaces with distinct edges or sharp features, which are often the most critical for defect detection in industrial inspection tasks. We demonstrate 0.5 m/s scanning speeds with MAE below 100 µm (11x faster than prior methods). A multi-reference Bayesian fusion strategy reduces MAE by 25.2% (vs. EMVS) and mitigates curvature errors. We also validate high-speed Braille reading 2.6x faster than previous approaches.
@@ I Introduction
Surface metrology and surface inspection are crucial elements in quality assurance across diverse industries, particularly aerospace and automotive manufacturing. Precise inspection is required to identify characteristics like paint quality, coating integrity, and subtle defects such as cracks, nicks, and dents [futterlieb2017air, yusra, Yuan2017]. Often, achieving a resolution of 0.1 mm or lower is necessary to accurately classify these features and ensure component integrity and safety [Agarwal2023]. Traditional contact-based methods, including high-precision profilometers [sekatskii2019analysis, Villarrubia1997] or microscopic techniques [Russell2001, Yang2018, Jonkman2020], offer high resolution locally but become exceedingly time-consuming when applied to large surface areas due to their sequential, point-by-point or small-patch measurement nature. Non-contact optical methods, such as cameras, laser scanners, or structured light systems [yusra, Salah2023, deng2024review, laser, zhang2018high, ramadan_vision-guided_2024], can significantly accelerate inspection by capturing data over wider areas. However, these methods often lack robustness; their performance can be compromised by variations in ambient lighting, motion blur when attempting high-speed scanning, or challenging surface optical properties like high reflectivity or transparency [Hausler2011]. This highlights a critical need for sensing modalities that combine high resolution and robustness with the capability for rapid inspection of large, potentially complex surfaces. Capturing the full 3D geometry is particularly important, as it enables the quantification of volumetric defects (e.g., dent depth and volume) and allows for direct comparison against a target CAD model, a critical step in modern quality assurance.
Vision-Based Tactile Sensors (VBTSs), particularly those derived from the GelSight concept [Yuan2017, Johnson2011, li2020review], have emerged as a powerful contact-based alternative offering a unique blend of high-resolution geometric and textural sensing. These sensors typically employ a camera to view the deformation of a soft, reflective elastomer membrane pressed against a target surface. This principle allows for detailed surface capture, often with inherent robustness to external lighting conditions due to internal illumination [Yuan2017, Dong2017]. Numerous variations of GelSight have been developed, enhancing capabilities for robotic manipulation, complex shape capture, and integration into compact form factors [Donlon2018, Liu2022, Zhao2023, Jing2025].
Despite their advantages for detailed local analysis, the standard planar configuration of most VBTSs necessitates a discrete ”press-and-lift” interaction mode. This process, where the sensor is repeatedly pressed, lifted, and moved, becomes prohibitively time-consuming when inspecting large surface areas, requiring numerous measurements to be stitched together [Agarwal2023], thereby limiting their use for efficient large-scale tactile measurement.
@@
Neuromorphic vision sensors, or event cameras, provide a promising technological solution to overcome the speed and motion blur limitations of conventional VBTSs [Gallego2022, Huang2022]. Unlike frame-based cameras, event cameras feature pixels that respond independently and asynchronously to logarithmic brightness changes, generating data with microsecond-level temporal resolution, high dynamic range, and inherent resistance to motion blur [Brandli2014, Posch2011, Lichtsteiner2008]. These characteristics are ideally suited for capturing the rapid, dynamic deformations occurring at the contact interface of a high-speed tactile sensor. Crucially, integrating event cameras into tactile sensing holds considerable potential for achieving high-resolution continuous tactile sensing at higher speeds with significantly lower computational power, particularly for surfaces with sparse details, compared to conventional VBTSs. While Neuromorphic VBTSs (NVBTS) have been explored for applications like robotic grasping and slip detection (e.g., Evetac [Funk2024]) or contact analysis [Ward-Cherrier2020, Sajwani2023, Faris2023, neutac_salah_2024], their potential for enabling high-speed, continuous 3D large surface reconstruction has not been realized.
To enable continuous tactile measurement, various designs moving beyond the static planar configuration have emerged. Attempts at direct sliding of planar sensors face significant friction, wear, and motion blur challenges [potdar_high-speed_2024]. Roller-based [Cao2023, Shimonomura2021] and belt-based [mirzaee_gelbelt_2025] VBTS designs replace the static contact with a dynamic rolling interface, allowing continuous data acquisition. TouchRoller, for example, demonstrated 2D texture mapping over moderate areas [Cao2023], while GelBelt provides a larger contact patch via its belt mechanism [mirzaee_gelbelt_2025]. Despite these mechanical advancements for measurement continuity, these designs face limitations. Reconstructing accurate 3D geometry from the depth-variant information gathered by cylindrical rollers is complex, primarily because the contact depth is not uniform across the curved sensing area [Li2023]. More critically, the operational speed of systems relying on conventional cameras is constrained. Frame rate limitations and motion blur become significant issues as scanning speed increases, particularly for VBTSs that require close-up imaging with a shallow depth of field to resolve fine surface details. Consequently, current roller/belt tactile sensors are restricted to relatively low velocities (e.g., typically well below 0.1 m/s for 3D reconstruction [Cao2023, mirzaee_gelbelt_2025]), insufficient for truly high-throughput industrial inspection of large parts. A significant gap exists for a tactile sensing system capable of combining the continuous scanning geometry of a roller with the high temporal fidelity and motion robustness of event-based vision for large-scale surface measurement.
@@
This paper introduces, to the best of our knowledge, the first high-speed neuromorphic vision-based roller tactile sensor specifically designed for large surface measurements (illustrated in Figure 1). The sensor is effective for 3D reconstruction of surfaces with distinct edges or sharp features, which are often the most critical for tactile feature recognition and defect detection. Our primary contributions are:
@@
    We introduce a novel rolling neuromorphic vision-based tactile sensor for 3D reconstruction and fine feature recognition of surfaces at speeds up to $0.5$ m/s, 11 times faster than previous continuous tactile sensing approaches. Furthermore, we rigorously evaluate its performance against the state-of-the-art in large surface inspection in terms of both speed and accuracy.
@@
    We adapt the event-based multi-view stereo (EMVS) method for curved tactile sensing by integrating a Bayesian Model Averaging (BMA) multi-reference depth fusion strategy, improving depth consistency and reducing curvature-induced errors. Moreover, a corresponding calibration routine is developed to optimize its parameters.
@@ II Related Works
Reconstructing a surface’s three-dimensional geometry is crucial in many applications, from automated manufacturing to quality control. Mechanical profilometers have long been used for high-precision surface measurements, meticulously tracing the surface point-by-point. However, this inherent sequential nature makes them slow, especially when dealing with complex geometries or large areas [Villarrubia1997]. Non-contact optical techniques, like laser scanning and structured light, offer a significant speed advantage by rapidly capturing surface data over a wider area [Salvi2004, Sansoni2009]. These methods boast resolutions around 0.1 mm [Johnson2011], but their reliance on light makes them susceptible to issues with highly reflective or transparent materials, as well as variations in ambient lighting conditions [Hausler2011]. Furthermore, certain laser scanning systems, particularly higher-power ones used for speed or range, can introduce eye-safety concerns requiring adherence to safety standards and protocols [IEC60825].
@@
For applications demanding the utmost precision, techniques like white light interferometry, confocal microscopy, and atomic force microscopy (AFM) are utilized [Russell2001, Yang2018, Jonkman2020]. These methods can resolve surface details down to the nanometer scale, revealing intricate textures and features. However, they often involve complex setups, lengthy measurement times, and significant expense, limiting their practical use for large-scale inspection tasks.
II-B Vision-Based and Neuromorphic Tactile Sensing
@@
Vision-Based Tactile Sensors (VBTSs) have emerged as compelling alternatives for surface sensing, offering robustness, versatility, and ease of use. The GelSight sensor pioneered this field, making use of a soft, deformable surface with camera-based observation to extract detailed surface information utilizing photometric stereo [Johnson2011]. By analyzing the deformation of the elastomer, typically coated with a reflective layer, VBTSs can capture high-resolution data on surface geometry and texture, even under varying lighting [Yuan2017, Dong2017]. From the original GelSight concept, a diverse range of VBTS designs have emerged to enhance performance and broaden application: GelSlim for more compact integrations [Donlon2018], GelSight Fin Ray for soft compliant grippers [Liu2022], and GelSight Svelte for human shape mimicking [Zhao2023], among others. Simulation tools have also evolved in parallel, accelerating the design and deployment of novel VBTS solutions [Agarwal2021, Wang2022, Akinola2025, Zhong2025].
@@
Despite these advantages, conventional VBTSs face challenges for continuous large-area scanning. The rigid backing of the elastomer limits its ability to slide smoothly across surfaces, leading to distortions and potential damage to the sensing membrane. As a result, most operate in a discrete “press-and-lift” mode, which is slow for large-area inspection; for example, the commercial GelSight Mobile™ Series 2 requires approximately 0.1 seconds of exposure time per measurement patch [GelSightMobile2024], making large-area coverage time-consuming [Agarwal2023]. Attempts at continuous sliding, such as with the DIGIT sensor for high-speed Braille reading [potdar_high-speed_2024], have shown potential but also revealed limitations: reliance on a conventional camera leads to motion blur at higher speeds, reducing recognition accuracy, while friction during sliding can cause rapid wear or damage to the elastomer. Protective layers can mitigate wear but may introduce optical diffusion, limiting fine texture resolution, and require precise contact control to ensure consistent data acquisition.
@@
Neuromorphic vision sensors, or event cameras, offer a transformative solution to these speed and motion blur limitations [Gallego2022]. Operating asynchronously at the pixel level, they detect brightness changes with microsecond temporal resolution, high dynamic range, low latency, and inherent suppression of motion blur [Brandli2014, Posch2011, Lichtsteiner2008]. These properties are well-suited to capturing the rapid, fine-scale elastomer deformations that occur during high-speed tactile scanning. Several Neuromorphic VBTSs (NVBTSs) have leveraged these advantages: Gaussian processes on event data for contact force estimation [Naeini2020], spiking neural networks for texture classification (NeuroTac) [Ward-Cherrier2020], event-based marker tracking for slip detection (Evetac) [Funk2024], graph neural networks for contact angle prediction (Tactigraph) [Sajwani2023], and event-based image representations for proprioception and contact analysis [Faris2023]. While these works validate the potential of NVBTSs, they have primarily focused on grasping-related tasks (slip, force, contact classification) or local texture analysis. Moreover, for these use cases, the algorithms used in combination with most NVBTS often result in decreased spatial resolution due to their reliance on tracking sparse markers, which is detrimental for surface reconstruction and depth estimation [Funk2024].
II-C Roller-Based and Large Surface Tactile Sensors
@@
Roller-based tactile sensors represent a promising avenue for achieving continuous tactile sensing, particularly for large surfaces where traditional planar sensors are inefficient. Early work explored cylindrical designs for specific detection tasks, such as the sensor developed by Shimonomura et al. [Shimonomura2021] for detecting foreign bodies like food impurities during rolling contact. Similarly, studies in robotic in-hand manipulation have successfully utilized a roller grasper, often equipped with high-resolution tactile feedback, for tasks like object insertion and regrasping [Lepert2023, Yuan2025], demonstrating the utility of rolling contact in dynamic scenarios beyond simple surface mapping.
@@
Transitioning towards surface property characterization, TouchRoller [Cao2023] demonstrated the potential for mapping surface textures by continuously rolling a sensor element across a fabric sample. It achieved coverage of an 8cm $\times$ 11cm area in 10 seconds, yielding a 2D texture map. Notably, its operational speed was limited to approximately 11 mm/s; the authors noted that higher speeds resulted in motion blur within the conventional camera frames, hindering performance. This motion blur problem is particularly acute for VBTSs, which require close-up imaging to resolve fine surface details. Operating at such close proximity means that even moderate rolling speeds translate into high velocities of features across the image sensor plane; during the camera’s finite exposure time, these features traverse a larger number of pixels, resulting in significant motion blur compared to imaging from a distance. Furthermore, TouchRoller did not perform 3D surface reconstruction. Addressing 3D reconstruction, GelRoller [zhang_gelroller_2024] employed a cylindrical design integrating photometric stereo techniques for local surface normal estimation from single frames. For reconstructing larger surfaces, they proposed stitching multiple local reconstructions using Iterative Closest Point (ICP) registration. While demonstrating 3D capability, this work did not explicitly analyze or optimize the operational speed of the rolling process.
@@
A fundamental challenge inherent to cylindrical roller designs, however, is that the tactile information gathered varies significantly with contact depth due to the sensor’s curvature. This leads to inconsistencies and potential distortions in the acquired tactile images, complicating accurate reconstruction. Researchers have explored methods to mitigate this depth inconsistency. Li et al. [Li2023] proposed a cyclic image fusion strategy, processing information from different contact phases within the rolling sequence using wavelet transforms to enhance information extraction and reduce bias from the curved contact.
@@
An alternative approach addresses the issue mechanically: the GelBelt sensor [mirzaee_gelbelt_2025] utilizes a flat elastomeric belt moving between two wheels, creating a larger, more uniform contact area akin to a conveyor. This design avoids the depth variation inherent in direct cylindrical contact and enabled continuous sensing at higher reported speeds, up to 45 mm/s. However, this design comes with some drawbacks. It relies on sliding motion to achieve high-speed operation, using an air gap and a smooth surface to reduce friction. While this allows GelBelt to move quickly, it creates challenges for lighting and can lead to wear over time, which may affect the system’s accuracy and repeatability. The design also struggles with curved surfaces. Its large, flat structure makes it hard to access concave areas, and it doesn’t perform well on convex surfaces either. In this case, the flat sensor doesn’t match the shape of the curved surface, leading to a limited sensing area, similar to problems seen in roller-based systems.
@@
Furthermore, despite these advances in measurement continuity, including GelBelt’s increased speed, systems relying on conventional cameras remain fundamentally constrained by the trade-off between speed and accuracy due to motion blur. As demonstrated by GelBelt’s own results showing decreased accuracy at higher speeds [mirzaee_gelbelt_2025], pushing scanning velocities further significantly degrades the quality of the captured tactile information and subsequent reconstructions. While naively increasing camera frame rates might seem like a straightforward solution to combat motion blur, it introduces substantial challenges regarding the increased computational complexity required for real-time processing, significantly larger data storage requirements, and higher energy consumption, potentially limiting practical deployment. In turn, the potential for using event-based vision in roller sensors to achieve similar resolution at low computational power and significantly higher speeds for continuous tactile sensing is considerable. However, this capability has yet to be demonstrated, particularly for demanding large-scale inspection tasks.
III Background - Event Camera Principles
Unlike conventional cameras that capture intensity information across all pixels at fixed frame rates, event cameras employ a fundamentally distinct, bio-inspired sensing paradigm [Gallego2022]. Each pixel operates independently and asynchronously, responding dynamically to changes in illumination. Specifically, a pixel monitors the logarithm of the incident light intensity, $L(x,y,t)=\log(I(x,y,t))$. An event $e_{k}=(x_{k},y_{k},t_{k},p_{k})$ is generated at coordinates $(x_{k},y_{k})$ only when the change in this logarithmic intensity, $\Delta L$, relative to its value at the time of the last event from that pixel, surpasses a predefined contrast threshold, $\pm C$. The generated event includes a high-resolution timestamp $t_{k}$ and a polarity $p_{k}\in\{-1,+1\}$, signifying the direction of the brightness change (decrease or increase, respectively). Following event generation, the pixel’s internal reference brightness level is updated, priming it for subsequent intensity variations. The output is thus not a sequence of frames, but a sparse, continuous stream of asynchronous events reflecting dynamic scene changes. The unique characteristics of event cameras offer substantial benefits for high-speed tactile inspection. Their inherent high temporal resolution, with events timestamped often at microsecond precision, allows for tactile sensing at high speeds without the motion blur experienced by other roller tactile sensors at similar speeds. Furthermore, event-based cameras benefit from low latency and high dynamic range allowing for potential reduction in computational load and robustness to varying illumination within the sensor housing respectively.
IV Methodology
IV-A Mechanical Design and Fabrication